# Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble

### v1

@@ Abstract
In this paper, we investigate the $n$-replica time evolution operator $\mathcal{U}_{n}(t)\equiv e^{\mathcal{L}_{n}t}$ for the Brownian Gaussian Unitary Ensemble (BGUE) using a graph-theoretic approach. We examine the moments of the generating operator $\mathcal{L}_{n}$, which governs the Euclidean time evolution within an auxiliary $D^{2n}$-dimensional Hilbert space, where $D$ represents the dimension of the Hilbert space for the original system. Explicit representations for the cases of $n=2$ and $n=3$ are derived, emphasizing the role of graph categorization in simplifying calculations. Furthermore, we present a general approach to streamline the calculation of time evolution for arbitrary $n$, supported by a detailed example of $n=4$. Our results demonstrate that the $n$-replica framework not only facilitates the evaluation of various observables but also provides valuable insights into the relationship between Brownian disordered systems and quantum information theory.
preprint: USTC-ICTS/PCFT-25-05
@@ 1 Introduction
Brownian disordered quantum systems have garnered significant interest across various fields, including many-body quantum chaos Chen_2019 ; S_nderhauf_2019 ; kalsi2024hierarchicalanalyticalapproachuniversal , quantum gravity Brady_2013 ; pelliconi2024gravitymesoscopic , and quantum information theory Brady_2013 ; Bentsen_2021 ; liu2023signaturescrambloneffectivefield ; Zhang_2023 ; Tang:2024kpv . Unlike systems with static (quenched) disorder, Brownian disorder is characterized by time-dependent randomness, allowing ensemble averaging to yield analytically tractable descriptions of non-equilibrium dynamics. A paradigmatic example is the Brownian Sachdev-Ye-Kitaev (SYK) model Saad:2018bqo , which is significantly trackable than the regular SYK Maldacena_2016 ; Maldacena:2016hyu . This has led to its widespread use as a toy model for quantum gravity Narovlansky:2023lfz ; Milekhin:2023bjv ; Stanford:2023npy ; Hayden_2007 , quantum error correction (QEC) Balasubramanian:2023xdp , out-of-time-ordered correlators (OTOCs) Sunderhauf:2019djv , and other fundamental aspects of quantum chaos Jian:2020krd ; Agarwal:2020yky ; Roberts_2017 ; Cotler_2017 ; jian2022lineargrowthcircuitcomplexity ; Peng:2022pfa ; fava2023designsfreeprobability .
Building on these developments, the recently proposed Brownian Gaussian Unitary Ensemble (BGUE) Tang:2024kpv offers an alternative setting to explore Brownian disorder in quantum systems. Unlike the standard Gaussian Unitary Ensemble (GUE), where the Hamiltonian is randomly drawn but remains fixed in time, BGUE describes a system evolving under a time-dependent noisy Hamiltonian,
$H_{ij}(t)=\eta_{ij}(t),\eta_{ij}(t)=\eta_{ij}^{\ast}(t)\in\mathbb{C},$ (1)
@@
where $\eta_{ij}(t)$ are independent Brownian Gaussian random variables with vanishing mean and nonzero variance
$\mathbb{E}\left(\eta_{ij}(t)\right)=0,\ \mathbb{E}\left(\eta_{ij}(t)\eta_{kl}(t^{\prime})\right)={J\over D}\delta_{il}\delta_{jk}\delta(t-t^{\prime})\ .$ (2)
The time evolution in this setting can be regarded as the limiting case of an infinitely long piecewise stochastic evolution
$U=e^{-\text{i}H_{1}t_{1}}e^{-\text{i}H_{2}t_{2}}\cdots e^{-\text{i}H_{n}t_{n}},$ (3)
where each $H_{i}$ is an independent realization of a GUE matrix. This model exhibits several remarkable properties, including rapid convergence to unitary $k$-designs dankert2005efficientsimulationrandomquantum ; Gross_2007 ; ambainis2007quantumtdesignstwiseindependence ; Dankert_2009 ; Cotler_2017 ; Roberts_2017 ; chen2024efficientunitarytdesignsrandom , hyperfast scrambling and the emergence of temperature in de Sitter space lin2022infinitetemperatureshot . Compared to conventional chaotic systems, the Brownian nature of BGUE ensures that ensemble-averaged quantities remain well-defined and analytically trackable, making it a powerful tool for exploring non-equilibrium dynamics.
One of the key motivations for studying BGUE lies in its connection to quantum information theory PhysRevADankert ; Gross2007 ; 4262758 ; Scott_2008 ; Roy_2009 ; Harrow_2009 ; Brand_o_2016 ; Nakata_2017 ; hunterjones2019unitarydesignsstatisticalmechanics . Random matrix models have long served as powerful tools for understanding quantum thermalization, entanglement growth, and operator spreading. In particular, BGUE has been shown to efficiently generate unitary designs, which play a central role in quantum information processing tasks such as classical shadow tomography Aaronson:2017qlb and quantum state reconstruction Tang:2024kpv . Furthermore, an essential feature of disordered quantum systems is their behavior under replica dynamics, a technique in which multiple copies of the system are introduced to probe statistical correlations. Many important quantum information measures—including entanglement entropy, operator norms, Rényi divergences, and multi-contour OTOC Haehl:2017pak —require analyzing multi-replica observables, making the study of $n$-replica time evolution in BGUE a crucial problem. However, existing studies have focused primarily on the cases $n=1,2$, leaving the general $n$-replica case largely unexplored. This serves as one of our motivations to study the general $n$-replica case and develop a systematic way to solve this question.
However, a fundamental challenge in extending the analysis to arbitrary $n$ is the rapid growth in computational complexity. The effective time evolution of the $n$-replica
@@
evolves within a Hilbert space of dimension $D^{2n}$, making direct matrix representations infeasible for large $n$. To simplify this question, we should note that the calculation of the $n$-replica time evolution operator $\mathcal{U}_{n}(t)$ can be reduced to finding the general expression for the moments of $\mathcal{L}_{n}$, allowing for a straightforward expansion:
@@
We find that the linearly independent terms, which we refer to as “graphs”, that appear in the moments of $\mathcal{L}_{n}$ are finite in number, allowing for a natural matrix representation. In the case of $n=2$, the author of Tang:2024kpv categorized 24 individual graphs into 8 groups, reducing the matrix dimension from 24 to 8. However, as $n$ increases, the number of individual graphs that emerge in the moments of $\mathcal{L}_{n}$ grows rapidly (see Table 1), necessitating a systematic method to group these graphs into appropriate categories to achieve a significantly smaller-dimensional representation.
In this work, we develop a graph-theoretic framework for analyzing the time evolution of $n$-replica observables in BGUE. Our approach generalizes and extends the graph-based categorization introduced for $n=2$, introducing a more structured notation that enables an algebraic treatment of replica dynamics. We demonstrate that the moments of $\mathcal{L}_{n}$ can be mapped onto a finite set of graph categories, significantly reducing computational complexity.
For $n=3$, we have 26 graph categories, so that $\mathcal{L}_{n}$ is a $26$-dimensional matrix. This enables us to obtain an explicit analytical expression for
$e^{\mathcal{L}_{3}t}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\ ,$ (6)
where the analytical expressions of $f_{a}(t)$ are listed in Appendix A. Furthermore, we discuss the generalization of our framework to arbitrary $n$ and present explicit results for $n=4$.
Compared to representation-theoretic approaches, our graph-based method offers a lower-dimensional irreducible representation, reducing numerical computation time by handling a lower-dimensional matrix. Previous approaches, such as those in Guo:2024zmr , relied on
@@
which allows for the determination of eigenvalues and wavefunctions for $\mathcal{L}_{n}$ by decomposing direct products of SU$(d)$ into irreducible representations. Moreover, these representations are constructed systematically using Young tableaux Hamermesh1962GroupTA . We will show the detailed result for $n=2$ at the end of Sec. 3.
The graph representation also presents a straightforward way to compute the expectation values of observables. In cases where symmetries like (7) do not exist, the methodology discussed here can serve as a valuable toolbox. Beyond its application to BGUE, our framework has broader implications to other Brownian disordered systems, like random unitary circuit Nahum_2017 ; Nahum_2018 ; von_Keyserlingk_2018 ; Zhou_2020 ; Fisher_2023 and random tensor networks Hayden_2016 ; Hosur:2015ylk . The techniques developed here could provide new insights into the replica method in many-body quantum chaos, as well as potential applications to problems in wormholes and quantum gravity Maldacena_2003 ; saad2019latetimecorrelationfunctions ; stanford2020quantumnoisewormholes ; Peng:2021vhs .
@@
The remainder of this paper is structured as follows. In Sec. 2, we introduce the notation and graph-based formalism for representing the time evolution operator $\mathcal{U}_{n}(t)$. In Sec. 3, we apply our framework to the case $n=2$ as a warm-up example, treated as a double check with the results from Tang:2024kpv . Sec. 4 extends this analysis to the non-trivial case $n=3$, where we show how to find the graph categories and assign the graphs into them. In Sec. 5, we present a general algorithm for arbitrary $n$ and illustrate its application to $n=4$. Finally, in Sec. 6, we discuss potential future directions and broader implications of our findings.
2 Analysis on Graphs
@@
At first, we introduce some notations for the graph representation of $\mathcal{U}_{n}$ and $\mathcal{L}_{n}$. We treat $\mathcal{U}_{n},\mathcal{L}_{n}$ as a $D^{2n}$-dimensional matrix, with row and column indices labeled by two $n$-length lists: $\{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}\}$ and $\{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}\}$. We then adapt this to the graph representation (taking $n=2$ as an example):
$\left(\mathcal{U}_{2}(t)\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $\equiv\mathbb{E}\left(U_{I_{1}I_{1}^{\prime}}U_{J_{1}J_{1}^{\prime}}^{*}U_{I_{2}I_{2}^{\prime}}U_{J_{2}J_{2}^{\prime}}^{*}\right)=\mathbb{E}\left[\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (8)
We label the indices in graphs using uppercase letters $I,J,I^{\prime},J^{\prime}$, where $I/I^{\prime}$ represent the left/right-hand side indices of a contour $H$, and $J/J^{\prime}$ denote the dual contour $H^{*}$. For simplicity, we label the contours (and dual contours) with lowercase letters $i,j,k,l$ (and their corresponding bars $\bar{i},\bar{j},\bar{k},\bar{l}$). Since we take $n=2$, we can also represent $\mathcal{U}_{2}$ as
$\mathcal{U}_{2}(t)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}$ (9)
@@
$\mathcal{L}_{n}=\mathsf{w}\mathbb{I}+{J\over D}\sum_{i\bar{j}}^{n}P_{i\bar{j}}-{J\over D}\sum_{ij}^{n}(X_{ij}+X_{\bar{i}\bar{j}})\quad,$ (10)
where $J$ is the strength of noise, $\mathbb{I}$ denotes identity operator. Other two kinds of operators are represented as below:
$P_{i\bar{j}}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$},~{}X_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\quad,$ (11)
where each line with two indices $i,j$ at the endpoints of the graph represents a “propagator” between the sites on the contour and the dual contour, i.e., $\delta_{ij}$. Other trivial propagators connecting sites on the same contour, which are represented as straight lines, are given by $\delta_{I_{i}I_{i}^{\prime}}$ or $\delta_{J_{i}J_{i}^{\prime}}$. Taking $n=2$ as an example, we have
$\left(P_{1\bar{2}}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}J_{2}}\delta_{I_{1}^{\prime}J_{2}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{I_{2}I_{2}^{\prime}}\ ,$
$\left(X_{12}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}I_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{J_{2}J_{2}^{\prime}}\ .$ (12)
And $\mathsf{w}$ in (10) is a factor that carries the indices of a graph
@@
Notice that the factor $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ commutes with all of the terms in $\mathcal{L}_{n}$. For instance,
@@
Since $\mathsf{w}=\mathsf{E}-nJ\mathbb{I}$, the factor $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ also commutes with all terms in $\mathcal{L}_{n}$.
2.1 Notations of a Graph
In this work, we employ a concise notation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ to represent an arbitrary graph. Here, $0\leq p\leq n$ denotes the number of paired vertices, and $\sigma\equiv\sigma_{H}\otimes\sigma_{H^{*}}\in S_{n}\otimes S_{n}$ encodes the permutation acting on the indices appearing on the right-hand side of the graph. The element of a $D^{2n}$-dimensional matrix $\mathsf{F}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}$ would be:

### v2

@@ Abstract
In this paper, we investigate the $n$-replica time evolution operator $\mathcal{U}_{n}(t)\equiv e^{\mathcal{L}_{n}t}$ for the Brownian Gaussian Unitary Ensemble (BGUE) through a graph-theoretic approach. We analyze the moments of the generator operator $\mathcal{L}_{n}$, which governs the Euclidean time evolution within an auxiliary $D^{2n}$-dimensional Hilbert space, where $D$ represents the dimension of the Hilbert space for the original system. Explicit representations for the cases of $n=2$ and $n=3$ are derived, highlighting the significance of graph categorization in simplifying complex calculations. Furthermore, we discuss a general formulation applicable to arbitrary $n$ and provide detailed examples, including the case of $n=4$. Our findings reveal that the $n$-replica framework not only facilitates the evaluation of various observables but also offers insights into the interplay between Brownian disordered systems and quantum information theory.
Keywords:
Brownian, GUE, Graph Representation
@@ 1 Introduction
Motivation
Brownian (disordered) systems hold significant importance in various fields, including quantum gravity and quantum information. Unlike the quenched disorder found in typical disordered systems, Brownian disorder results in a local theory after taking the ensemble average, making these systems considerably easier to solve. For example, the Brownian SYK model, a crucial toy model in quantum gravity, is much more manageable than its regular counterpart Maldacena_2016 ; Maldacena:2016hyu . Furthermore, the results obtained from the Brownian SYK can be utilized to evaluate the spectral form factor (SFF) of the standard SYK model Saad:2018bqo . The connection between the Brownian SYK and de Sitter spacetime has also been explored Narovlansky:2023lfz ; Milekhin:2023bjv ; Stanford:2023npy .
Additionally, Brownian systems can be employed to achieve unitary $k$-designs dankert2005efficientsimulationrandomquantum ; Gross_2007 ; ambainis2007quantumtdesignstwiseindependence ; Dankert_2009 , which are distributions of unitaries that reproduce the first $k$ moments of the Haar measure. Unitary designs play a prominent role in various areas of quantum information theory PhysRevADankert ; Gross2007 ; 4262758 ; Scott_2008 ; Roy_2009 ; Harrow_2009 ; Brand_o_2016 ; Nakata_2017 ; hunterjones2019unitarydesignsstatisticalmechanics , many-body quantum chaos Roberts_2017 ; Cotler_2017 ; jian2022lineargrowthcircuitcomplexity ; fava2023designsfreeprobability , and models of black holes Hayden_2007 .
Recently, a model known as the Brownian Gaussian Unitary Ensemble (BGUE) was explored in Tang:2024kpv , alongside other similar models discussed in Guo:2024zmr . These studies consider a closed quantum system characterized by a noisy Hamiltonian
$H_{ij}(t)=\eta_{ij}(t),\eta_{ij}(t)=\eta_{ji}^{*}(t)\in\mathbb{C}$ (1)
@@
where $1\leq i,j\leq D$. $D$ is the dimension of the Hilbert space. And $\eta_{ij}(t)$ are independent Brownian Gaussian random variables with vanishing mean and non-vanishing variances
$\mathbb{E}\left(\eta_{ij}(t)\right)=0,~\mathbb{E}\left(\eta_{ij}(t)\eta_{kl}(t^{\prime})\right)={J\over D}\delta_{il}\delta_{jk}\delta(t-t^{\prime})\ .$ (2)
Here, $\mathbb{E}$ denotes the disorder average. The model can be regarded as an infinite version of the evolution $U=e^{-\text{i}H_{1}t_{1}}e^{-\text{i}H_{2}t_{2}}\cdots e^{-\text{i}H_{n}t_{n}},\text{i}=\sqrt{-1}$, where each $H_{i}$ is independently drawn from the Gaussian Unitary Ensemble (GUE). Researchers consider this piecewise evolution and expect it to approach the Haar measure much faster than the pure evolution $e^{-iHt}$ Cotler_2017 ; Roberts_2017 ; chen2024efficientunitarytdesignsrandom on a timescale independent of $D$. Similar to the Brownian SYK model, it was found that the BGUE exhibits hyperfast scrambling and the emergence of temperature lin2022infinitetemperatureshot . In contrast to systems without disorder, when dealing with general $n$-replica observables in this model, we ultimately obtain an effective time evolution on $2n$-contours:
@@
where $U(t)\equiv e^{-\text{i}Ht}$, $\mathcal{L}_{n}$ is the generator operator of $\mathcal{U}_{n}(t)$, functioning as the “Hamiltonian” for the Euclidean time evolution in the $D^{2n}$-dimensional Hilbert space $\mathcal{H}^{\otimes 2n}$. The calculation method bears similarities to random tensor network models Hayden_2016 and random unitary circuit models Nahum_2017 ; Nahum_2018 ; von_Keyserlingk_2018 ; Zhou_2020 ; Fisher_2023 . In Tang:2024kpv , the author has derived explicit analytical results for one-replica ($n=1$) and two-replica ($n=2$) time evolutions, enabling the evaluation of two-point functions, OTOCs, entropy, and other quantities in the model. For instance, we find a non-decaying two-point function as $t\to\infty$ Maldacena_2003 ; saad2019latetimecorrelationfunctions and non-vanishing fluctuations of the two-point function with a zero mean value stanford2020quantumnoisewormholes . Moreover, the application of the model to $n$-design and classical shadow tomography is also discussed. As argued in the paper, many interesting observables require three (or more) replicas, such as the max-version shadow norm. This serves as one of our motivations to study the general $n$-replica case. Note that the calculation of the $n$-replica time evolution operator $\mathcal{U}_{n}(t)$ can be reduced to finding the general expression for the moments of $\mathcal{L}_{n}$, allowing for a straightforward construction of $\mathcal{U}_{n}(t)$
@@
We find that the linearly independent terms, which we refer to as “graphs,” that appear in the moments of $\mathcal{L}_{n}$ are finite in number, allowing for a natural matrix representation. For the case of $n=2$, author in Tang:2024kpv previously addressed this problem by analyzing the features of the 24 individual graphs and categorizing them into 8 groups, effectively reducing the dimension of the matrix representation from 24 to 8. However, as $n$ increases, the number of individual graphs that emerge in the moments of $\mathcal{L}_{n}$ grows rapidly (see Table 1), creating the challenge of how to effectively group these graphs into appropriate categories to achieve a significantly smaller-dimensional matrix representation.
In this paper, we introduce a systematic notation for the graphs in the model, which facilitates a purely algebraic approach to the graph classification problem. This method enables us to derive a compact matrix representation for $\mathcal{L}_{n}$, making it a valuable tool for investigating the complex behaviors governed by this operator. We anticipate that the methodology presented in this paper can be applied to other models involving Brownian disorder.
The model
Here, we can generalize the BGUE by introducing an intrinsic spectral term
$H_{ij}=E_{i}\delta_{ij}+\eta_{ij}(t)\ .$ (5)
Summary of results
In this work, we establish a general framework to address the $n$-replica time evolution problem. We provide detailed calculations for the cases of $n=2$ and $n=3$ as illustrative examples, presenting the analytical expressions for the corresponding evolution operators $\mathcal{U}_{2}$ and $\mathcal{U}_{3}$. For $n=3$, we have 26 graph categories, so that $\mathcal{L}_{n}$ is a $26$-dimensional matrix, then we obtain $e^{\mathcal{L}_{3}t}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$, the analytical expressions of $f_{a}(t)$ are listed in Appendix A.
Note Added
After completing the paper, we found an alternative approach discussed in Guo:2024zmr . Noticing
@@
which allows for the determination of eigenvalues and wavefunctions for $\mathcal{L}_{n}$ by decomposing direct products of SU(d)(d) into irreducible representations. Moreover, these representations are constructed systematically using Young tableaux Hamermesh1962GroupTA . Thus, they have provided a systematic method for calculating any observable in the Brownian GUE. [^1] Similarly, one can solve Brownian GOE and Brownian GSE. We believe that the method presented in this paper is fundamentally similar to that in Guo:2024zmr , albeit expressed in a different form. In cases where symmetries like Eq.(6) do not exist, the methodology discussed here can serve as a valuable toolbox.
@@
We first introduce the denotation for graphs in Sec. 2. Then we study the $n=2$ case as a warm-up in Sec. 3. Next, we evaluate the non-trivial $n=3$ case in Sec. 4. Finally, we present a universal algorithm for the general $n$ case in Sec. 5, where the $n=4$ case is also discussed. Lastly, we provide some comments on future research in Sec. 6.
2 Denotation for graphs
@@
At first, we introduce some notations for the graph representation of $\mathcal{U}_{n}$ and $\mathcal{L}_{n}$. We treat $\mathcal{U}_{n}/\mathcal{L}_{n}$ as a $D^{2n}$-dimensional matrix, with row and column indices labeled by two $n$-length lists: $\{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}\}$ and $\{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}\}$. We then adapt this to the graph representation (taking $n=2$ as an example):
$\left(\mathcal{U}_{2}(t)\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $\equiv\mathbb{E}\left(U_{I_{1}I_{1}^{\prime}}U_{J_{1}J_{1}^{\prime}}^{*}U_{I_{2}I_{2}^{\prime}}U_{J_{2}J_{2}^{\prime}}^{*}\right)=\mathbb{E}\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (7)
We label the indices in graphs using uppercase letters $I,J,I^{\prime},J^{\prime}$, where $I/I^{\prime}$ represent the left/right-hand side indices of a contour $H$, and $J/J^{\prime}$ denote the dual contour $H^{*}$. For simplicity, we label the contours (and dual contours) with lowercase letters $i,j,k,l$ (and their corresponding bars $\bar{i},\bar{j},\bar{k},\bar{l}$). Therefore, we also denote
$\mathcal{U}_{2}(t)=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}$ (8)
@@
$\mathcal{L}_{n}=\mathsf{w}\mathbb{I}+{J\over D}\sum_{i\bar{j}}^{n}P_{i\bar{j}}-{J\over D}\sum_{ij}^{n}(X_{ij}+X_{\bar{i}\bar{j}})\ .$ (9)
Here, $\mathbb{I}$ denotes the identity operator, and we sometimes simply denote it as $I$. The other two kinds of operators are represented below:
$P_{i\bar{j}}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$},~X_{ij}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (10)
where each line with two indices $i,j$ at the endpoints of the graph represents a ”propagator,” i.e., $\delta_{ij}$. Other trivial contours, represented by straight lines, are given by $\delta_{I_{i}I_{i}^{\prime}}$ or $\delta_{J_{i}J_{i}^{\prime}}$. Taking $n=2$ as an example, we have
$\left(P_{1\bar{2}}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $=\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}J_{2}}\delta_{I_{1}^{\prime}J_{2}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{I_{2}I_{2}^{\prime}}\ ,$
$\left(X_{12}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $=\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}I_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{J_{2}J_{2}^{\prime}}\ .$ (11)
And $\mathsf{w}$ in Eq. (9) is a factor that carries the indices of a graph
@@
Notice that the factor $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ commutes with all of the terms in $\mathcal{L}_{n}$, e.g.,
@@
Since $\mathsf{w}=\mathsf{E}-nJ\mathbf{1}$, the factor $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ also commutes with all terms in $\mathcal{L}_{n}$.
2.1 Notation for a graph
In this work, we employ the concise notation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ to represent an arbitrary graph. Here, $0\leq p\leq n$ denotes the number of paired vertices, and $\sigma\equiv\sigma_{1}\otimes\sigma_{2}\in S_{n}\otimes S_{n}$ encodes the permutation acting on the indices appearing on the right-hand side of the graph (or equivalently, a $D^{2n}$-dimensional matrix $\mathsf{F}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}$: